# RGB-Th-Bench: A Dense benchmark for Visual-Thermal Understanding of Vision Language Models

### v1

@@ 1 Introduction
Vision-Language Models (VLMs) [8, 52, 53, 12, 57, 39, 38, 32, 6, 47, 3, 44] have made remarkable progress in recent years, leveraging large-scale image-text datasets to align visual representations with the powerful reasoning capabilities of Large Language Models (LLMs) [43, 49, 17, 7, 11]. To systematically evaluate their multimodal understanding, numerous benchmarks have been developed, such as MME [18], MMBench [40], SEED-Bench [27], and $M$-Vet [54]. These benchmarks have become widely recognized for assessing various vision-language skill dimensions, including core perception tasks, domain knowledge, reasoning abilities, and application-specific tasks.
@@ 2 Related Work
LLVIP [24] 15,488 Yes No $L$ Ped Det Peds High-quality alignment
@@
FLIR [4] 26,442 No No Self-Driving, Obj Det 15 Ctg. $B$ ann. provided
@@
Recent Benchmarks for Vision-Language Models. To systematically assess the capabilities of VLMs, several benchmarks have been introduced, such as MME [18], MMBench [40], SEED-Bench [27], MMMU [56], $M$-Vet [54], and NaturalBench [29]. These benchmarks have become widely recognized as standard evaluation tools, as reflected in their prominence on performance leaderboards [14].
Despite their broad adoption, these benchmarks primarily assess single-image text comprehension and lack support for tasks involving multiple images or interleaved image-text sequences. While this was a reasonable limitation in earlier VLMs designed for single-image inputs, recent advancements [57, 52, 53, 43, 44] have enabled models to process multiple images, sequential video frames, and arbitrarily interleaved multimodal inputs. In response to this shift, new benchmarks such as $M$-Vet2 [55], SeedBench2 [26], and Seed-Bench2-Plus [28] have been introduced as extensions to their predecessors.
@@ 3.2 Evaluation Dimensions
For each sample pair, there are a total of 56 questions across 14 core evaluation or skill dimensions, resulting in four independent Yes/No questions per skill dimension for each sample. This way, we make sure that every sample pair is thoroughly tested from all possible skill dimensions to make our benchmark a dense evaluation setting, which is unlike the benchmarks that ask only one or two questions from each sample image, like MME [18], $M$-Vet [54] and Seed-Bench2 [26]. These skills are grouped into two types of prompt:

### v3

@@ 1 Introduction
Vision-Language Models (VLMs) [8, 52, 53, 12, 57, 39, 38, 32, 6, 47, 3, 44] have made remarkable progress in recent years, leveraging large-scale image-text datasets to align visual representations with the powerful reasoning capabilities of Large Language Models (LLMs) [43, 49, 17, 7, 11]. To systematically evaluate their multimodal understanding, numerous benchmarks have been developed, such as MME [18], MMBench [40], SEED-Bench [27], and MM-Vet [54]. These benchmarks have become widely recognized for assessing various vision-language skill dimensions, including core perception tasks, domain knowledge, reasoning abilities, and application-specific tasks.
@@ 2 Related Work
LLVIP [24] 15,488 Yes No LL Ped Det Peds High-quality alignment
@@
FLIR [4] 26,442 No No Self-Driving, Obj Det 15 Ctg. BB ann. provided
@@
Recent Benchmarks for Vision-Language Models. To systematically assess the capabilities of VLMs, several benchmarks have been introduced, such as MME [18], MMBench [40], SEED-Bench [27], MMMU [56], MM-Vet [54], and NaturalBench [29]. These benchmarks have become widely recognized as standard evaluation tools, as reflected in their prominence on performance leaderboards [14].
Despite their broad adoption, these benchmarks primarily assess single-image text comprehension and lack support for tasks involving multiple images or interleaved image-text sequences. While this was a reasonable limitation in earlier VLMs designed for single-image inputs, recent advancements [57, 52, 53, 43, 44] have enabled models to process multiple images, sequential video frames, and arbitrarily interleaved multimodal inputs. In response to this shift, new benchmarks such as MM-Vet2 [55], SeedBench2 [26], and Seed-Bench2-Plus [28] have been introduced as extensions to their predecessors.
@@ 3.2 Evaluation Dimensions
For each sample pair, there are a total of 56 questions across 14 core evaluation or skill dimensions, resulting in four independent Yes/No questions per skill dimension for each sample. This way, we make sure that every sample pair is thoroughly tested from all possible skill dimensions to make our benchmark a dense evaluation setting, which is unlike the benchmarks that ask only one or two questions from each sample image, like MME [18], MM-Vet [54] and Seed-Bench2 [26]. These skills are grouped into two types of prompt: